OpenStudy (anonymous):

A glass vase is shaped like a cylinder and has a radius of 2.5 inches. The vase holds 170 cubic inches of water. How tall is the vase?

OpenStudy (anonymous):

@roz_mar23 @tester97

OpenStudy (anonymous):

H=V/3.14R^2 H=170/3.14*2.5^2=8.658 inches

OpenStudy (anonymous):

oooh i see now! thanks again! you're a lifesaverr!!

OpenStudy (anonymous):

Ur welcome

OpenStudy (anonymous):

I substituted in the next question into the formula and i got it incorrect. I'm really not great at math..I just don't know what I'm doing wrong. @roz_mar23

OpenStudy (anonymous):

What was the Q?

OpenStudy (anonymous):

If a cone-shaped vase can hold 45 cubic centimeters of water and has a radius of 2 inches, what is the height of the vase?

OpenStudy (anonymous):

H=3*V/3.14R^2 H=3*45/3.14*2^2=10.7429inches This is a cone not a cylinder

OpenStudy (anonymous):

oh my goodness...Facepalm** thanks again!

OpenStudy (anonymous):

Ur welcome